OpenStudy (anonymous):

Find the radian measure of an angle of -280 degrees

OpenStudy (anonymous):

We need to use 1 degree = pi/180 radians getting this ???

OpenStudy (anonymous):

Yeah I get that formula.

OpenStudy (anonymous):

Now multiply by -280 to both sides of equation 1 degree = pi/180 radians can you try ??

OpenStudy (anonymous):

1 Degree*-280=pi/180*-280 -280 degrees=pi/180*-280

OpenStudy (anonymous):

correct.....

OpenStudy (anonymous):

Now we need to reduce fraction can you try ?

OpenStudy (anonymous):